The label on the Listerine Cool Mint Antiseptic mouthwash says that it is 21.6% alcohol by volume. The label on the Scope Original Mint mouthwash says that it is 15.0% alcohol by volume. If Hans mixes 6 ounces of the Listerine with 4 ounces of the Scope, what is the percent alcohol content of the mixtures?
Answer by stanbon(75887)

The label on the Listerine Cool Mint Antiseptic mouthwash says that it is 21.6% alcohol by volume. The label on the Scope Original Mint mouthwash says that it is 15.0% alcohol by volume. If Hans mixes 6 ounces of the Listerine with 4 ounces of the Scope, what is the percent alcohol content of the mixtures?
-----------
Equation:
alcohol + alcohol = alcohol
0.216*6 + 0.15*4 = x(6+4)
------
Multiply thru by 1000 to get:
216*6 + 150*4 = 10000x
---
x = 18.96%
===============
Chjeers,
Stan H.